Describe how surge tower can help in reducing pressure in pipeline subjected to water hammer phenomena. A surge tower is required to be installed at the downstream end of a tunnel in a hydroelectric scheme to dampen the effect of waterhammer in the case of pump or electrical failures when the flow to the turbine is stopped suddenly. It has been anticipated that the diameter of the surge tower to be 10 m. Determine the height required for the surge tower. The surge tower is not to overflow. Calculate also the time required for the water level in the surge tank to reach its maximum value. The friction and minor losses can be neglected. The information of the tunnel is as follows: Length of tunnel = 10000 m Diameter of tunnel = 3 m Steady flow in the tunnel = 36 m3 /s
Answer #1

surge tank is used to nutralise the pressure in the penstok.when the valve closes partially to reduce flow there will be sudden rise in pressure due to excess water , but when surge tank is used the excess water flow to surge tank and pressure is maintained.

when valve is opened to increase flow , that moment the water in penstok is in short to meet the increased demand and there is a reduction of pressure. but when surge tank is used, water in the surge tank flows out to meet the increased demand and thus maintains the pressure.

water hammering is caused by increase in pressure. surge tank reduces it as it maintains the pressure.height of tank required Zmar - Q Sa Asa At As g Area of tanke As = T/10* = 78.54 m² 2 area of tunnel At T/A =7.07 m² x 32 Y Q